OpenStudy (anonymous):

Help :-) f(g(x) question!! F(x)= 5x+6 G(x)= 1/5 (x-6) When I do f(g(x) and g(f(x) my answer comes out to X! Am I doing this wrong????

OpenStudy (anonymous):

Hey get this wait exactly 30 seconds

OpenStudy (noelgreco):

No, it simply indicates that the functions are inverses.

OpenStudy (anonymous):

Yeah thats true

OpenStudy (anonymous):

So would my answer simply be x for f(g(x? Thank you

OpenStudy (noelgreco):

Yes.

OpenStudy (anonymous):

No you're doing it wrong sorta right.

OpenStudy (anonymous):

Explain

OpenStudy (anonymous):

Please

OpenStudy (noelgreco):

Any time you take a function of the function's inverse, the solution will simply be the variable.

OpenStudy (anonymous):

Ohhh how so?

OpenStudy (anonymous):

It cancels the inverse out pretty much

OpenStudy (anonymous):

Give me an example

OpenStudy (noelgreco):

Yeah, you could say that. F(G(x)) = F(1/5(x-6)) = 5(1/5(x-6) + 6 = x - 6 +6 = x Same thing if you take G(F(x))

OpenStudy (anonymous):

Ohh...I see u taking Algebra too ?

OpenStudy (anonymous):

I mean Alegbra TWO?

OpenStudy (noelgreco):

me?

OpenStudy (anonymous):

Yeah u?

OpenStudy (noelgreco):

No, I'm a retired math teacher having fun.